# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit
$\langle sl,\textbf{R}|s^{\prime}l^{\prime},\textbf{R}^{\prime}\rangle&=\delta_{ss^{\prime}}\delta_{ll^{\prime}}\delta_{\textbf{R}\textbf{R}^{\prime}},\\ \langle sl,\textbf{k}|s^{\prime}l^{\prime},\textbf{k}^{\prime}\rangle&=N_{\rm cell}\delta_{ss^{\prime}}\delta_{ll^{\prime}}\delta_{\textbf{k}\textbf{k}^{\prime}},\\ \langle\psi_{n\textbf{k}}|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle&=N_{\rm cell}\delta_{nn^{\prime}}\delta_{\textbf{k}\textbf{k}^{\prime}}\,.\end{aligned}\right.$ (20)
So, from Eq. (16), we get
$\epsilon_{\gamma\alpha\beta}$ ${\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P\}=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\langle\partial_{k_{x}}u_{n\textbf{k}}|H_{\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ (21)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ}}\sum_{m\in{\rm emp}}\sum_{s\,l}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})E_{m\textbf{k}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle u_{m\textbf{k}}|\partial_{k_{\beta}}u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}.$
Similarly, Eq. (17) becomes
$\epsilon_{\gamma\alpha\beta}$ ${\rm Tr}\{Qr_{\alpha}PHPr_{\beta}Q\}=-\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}$ (22)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}\sum_{s\,l}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}.$
@@
$M_{z}$ $=\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ}}\sum_{m\in{\rm emp}}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})E_{m\textbf{k}}\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{m\textbf{k}}|u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$ (23)
$-\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$
@@
where the layer-resolved orbital magnetization of the $l$-th layer $M_{z}^{(l)}$ is defined by summing all the degrees of freedom except layer $l$. As discussed before, when considering Chern insulators having surface contribution, one needs to replace $H\rightarrow H-\mu$ and $E_{n\textbf{k}}\rightarrow E_{n\textbf{k}}-\mu$ in the previous formula. Then, Eq. (3) of the main text is then derived.
@@ I.3 Layer Chern number
$C_{\gamma}(\textbf{r})=\frac{i}{4\pi}\epsilon_{\gamma\alpha\beta}\langle\textbf{r}|Pr_{\alpha}Qr_{\beta}P-Qr_{\alpha}Pr_{\beta}Q|\textbf{r}\rangle.$ (24)
@@
II II. Layer-resolved orbital magnetization under finite vertical electric field
In this section, we show that $M_{z}^{(l)}$ varies linearly with weak electric field for both types of stacking, independent of number of layers. Here, we calculate $M_{z}^{(l)}$ using the definition Eq. (23).
Figure 1: Layer-resolved $M_{z}^{(l)}$ of the slab under several $\mathcal{E}_{z}$ for different stacking ways and total layer numbers: (a) AA stacking, 5 layers; (b) AA stacking, 6 layers; (c) AB stacking, 5 layers; (d) AB stacking, 6 layers.
III III. Perturbation theory for layer-resolved orbital magnetoelectric response
When a vertical electric field is applied to a slab, our perturbed Hamiltonian reads: $H=H_{0}+e\mathcal{E}z$. Suppose the electric field is so weak such that bulk gap is not closed and the potential drop between the top and bottom layer is the smallest energy scale in the system. The idea is to express Eq. (23) of perturbed Hamiltonian $H$ using the information of $H_{0}$. By singling out the terms proportional to $\mathcal{E}$, the layer-resolved orbital magnetoelectric response $\alpha_{zz}^{(l)}$ is then derived by $\partial M_{z}^{(l)}/\partial\mathcal{E}$.
Let us first focus on the first-order correction in $\mathcal{E}$ to all terms in Eq. (23). First, the correction in wavefunction reads
$|u_{n\textbf{k}}\rangle$ $=|u_{n\textbf{k}}^{(0)}\rangle+\sum_{n^{\prime}\neq n}\frac{\frac{1}{N_{\rm cell}}\langle u_{n^{\prime}\textbf{k}}^{(0)}|e\mathcal{E}z|u_{n\textbf{k}}^{(0)}\rangle}{E_{n\textbf{k}}^{(0)}-E_{n^{\prime}\textbf{k}}^{(0)}}|u_{n^{\prime}\textbf{k}}^{(0)}\rangle$ (26)
@@
$C_{sl,n}(\textbf{k})$ $\equiv\langle sl,\textbf{k}|u_{n\textbf{k}}\rangle$ (27)
@@
$|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle=-iA_{n,\alpha}(\textbf{k})|u_{n\textbf{k}}\rangle+\sum_{m\neq n}\frac{\frac{1}{N_{\rm cell}}|u_{m\textbf{k}}\rangle\langle u_{m\textbf{k}}|}{E_{n\textbf{k}}-E_{m\textbf{k}}}\,(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{n\textbf{k}}\rangle,$ (28)
with Berry connection $A_{n,\alpha}(\textbf{k})=i\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle$. Note from Eq. (23) that we only need to calculate $\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ with $m\neq n$ so we can omit the first term in Eq. (28). We also note that $\partial_{k_{\alpha}}H_{\textbf{k}}=\partial_{k_{\alpha}}H_{\textbf{k}}^{(0)}$ with $H_{\textbf{k}}^{(0)}=e^{-i\textbf{k}\cdot\textbf{r}}H_{0}\,e^{i\textbf{k}\cdot\textbf{r}}$. Therefore we need to single out the first-order correction to $|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle$ by expanding all other terms. The simplest one would be $E_{n\textbf{k}}$
$E_{n\textbf{k}}$ $=E_{n\textbf{k}}^{(0)}+\langle u_{n\textbf{k}}^{(0)}|e\mathcal{E}z|u_{n\textbf{k}}^{(0)}\rangle$ (29)
@@
Let’s now expand the second term in Eq. (28). Since
@@
$\frac{1}{N_{\rm cell}}\sum_{m\neq n}\frac{|u_{m\textbf{k}}\rangle\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}H_{\textbf{k}}|u_{n\textbf{k}}\rangle}{E_{n\textbf{k}}-E_{m\textbf{k}}}$ (30)
@@
With all the ingredients derived above, we can expand Eq. (23) to the first-order of $\mathcal{E}$ field. Typically, we need to treat terms like:
@@
((*)( ∗ ) $=C_{sl,n}^{(0)}(\textbf{k})C_{sl,n^{\prime}}^{(0)*}(\textbf{k})E_{m\textbf{k}}^{(0)}\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{m\textbf{k}}^{(0)}|u_{n^{\prime}\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ (31)
@@
$\langle u_{n\textbf{k}}^{(0)}|u_{m^{\prime}\textbf{k}}^{(1)}\rangle_{\Omega_{0}}$ $=\sum_{s\,l}\sum_{m^{\prime\prime}\neq m^{\prime}}\langle u_{n\textbf{k}}^{(0)}|u_{m^{\prime\prime}\textbf{k}}^{(0)}\rangle_{\Omega_{0}}\frac{C_{sl,m^{\prime\prime}}^{(0)*}(\textbf{k})C_{sl,m^{\prime}}^{(0)}(\textbf{k})}{E_{m^{\prime}\textbf{k}}^{(0)}-E_{m^{\prime\prime}\textbf{k}}^{(0)}}\,e\mathcal{E}ld$ (35)
@@
$v_{\alpha,nm}(\textbf{k})$ $\equiv\langle u_{n\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}},$ (36)
@@
$\langle u_{n\textbf{k}}^{(1)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ $=\sum_{n^{\prime}\neq n}\sum_{s\,l}\frac{C_{sl,n^{\prime}}^{(0)}(\textbf{k})C_{sl,n}^{(0)*}(\textbf{k})}{E_{n\textbf{k}}^{(0)}-E_{n^{\prime}\textbf{k}}^{(0)}}\,e\mathcal{E}ld\langle u_{n^{\prime}\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ (37)
@@
$\langle u_{n\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(1)}\rangle_{\Omega_{0}}$ $=\sum_{m^{\prime}\neq m}\sum_{s\,l}\Big{(}\frac{D_{sl,mm^{\prime}}(\textbf{k})}{\Delta_{\textbf{k},mm^{\prime}}}\tilde{\mathcal{E}}\,lv_{\alpha,m^{\prime}n}(\textbf{k})\Big{)}^{*}$ (38)
@@
So, Eq.(34) becomes:
$\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(1)}\rangle_{\Omega_{0}}=$ $-\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}^{2}}\sum_{s\,l}\big{(}D_{sl,mm}(\textbf{k})-D_{sl,nn}(\textbf{k})\big{)}\tilde{\mathcal{E}}\,l$ (39)
@@
Eq.(33) becomes:
$\langle u_{n\textbf{k}}^{(1)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}=\sum_{n^{\prime}\neq n,m;s\,l}\frac{D_{sl,nn^{\prime}}(\textbf{k})v_{\alpha,n^{\prime}m}(\textbf{k})}{\Delta_{\textbf{k},nn^{\prime}}\Delta_{\textbf{k},mn^{\prime}}}\,\tilde{\mathcal{E}}\,l.$ (40)
Eq.(32) becomes:
$\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}=\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}}.$ (41)
In Eq. (27), the first-order term becomes:
$C_{sl,n}^{(1)}(\textbf{k})=\sum_{n^{\prime}\neq n;s^{\prime}l^{\prime}}\frac{D_{s^{\prime}l^{\prime},n^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},nn^{\prime}}}\,\tilde{\mathcal{E}}\,l^{\prime}C_{sl,n^{\prime}}^{(0)}(\textbf{k}).$ (42)
@@
So, we decompose Eq. (31) into four contributions: $(B_{1})$ to $(B_{4})$.
We now explicit the first-order correction in $|u_{n\textbf{k}}\rangle$: Eq.(39) and Eq.(40)
$\langle u_{n\textbf{k}}^{(1)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle+\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(1)}\rangle\equiv-iA_{\alpha,nm}^{(1)}\qquad\qquad\quad({\rm correction\;in\;Berry\;connection})$ (43)
@@
Now, we are ready for Eq. (23) of the perturbed system. The first part reads
@@
Let us first look what to change if considering the correction in $\mu$. First, only substitutions $E_{m\textbf{k}}^{(0)}\rightarrow E_{m\textbf{k}}^{(0)}-\mu^{(0)}$ should be made in $(B_{2})$ and $(B_{4})$ because the rest of terms is already of first-order. As for $(B_{3})$, we should take $\mu^{(1)}$ into consideration. Our solution to $\mu^{(1)}$ is to use an auxillary system with OBC, conjugate to the system with PBC. Here, we mean “conjugate” by that both systems have the same level of discretization, in other words the back and forth Fourier transformation is normalized to unity. For example, if we consider a system with PBC of k mesh $100\times 100$ in the Brillouin zone, its conjugate OBC system is a superlattice of $100\times 100$ unit cells. For isolated system, a weak external electric field cannot change the particle number of system. So, if the system is half-filled, one determine $\mu^{(1)}$ by
$\mu^{(1)}=\frac{1}{2}(E^{(1)}_{{\rm OBC},v}+E^{(1)}_{{\rm OBC},c})$ (44)
@@
$\mu^{(1)}=\frac{1}{2}\sum_{\textbf{R},sl}(\tilde{C}_{sl,v}^{(0)\,*}(\textbf{R})\tilde{C}_{sl,v}^{(0)}(\textbf{R})+\tilde{C}_{sl,c}^{(0)\,*}(\textbf{R})\tilde{C}_{sl,c}^{(0)}(\textbf{R}))\,\tilde{\mathcal{E}}l$ (45)

### v3

@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit: second approach
$\langle sl,\textbf{R}|s^{\prime}l^{\prime},\textbf{R}^{\prime}\rangle&=\delta_{ss^{\prime}}\delta_{ll^{\prime}}\delta_{\textbf{R}\textbf{R}^{\prime}},\\ \langle sl,\textbf{k}|s^{\prime}l^{\prime},\textbf{k}^{\prime}\rangle&=N_{\rm cell}\delta_{ss^{\prime}}\delta_{ll^{\prime}}\delta_{\textbf{k}\textbf{k}^{\prime}},\\ \langle\psi_{n\textbf{k}}|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle&=N_{\rm cell}\delta_{nn^{\prime}}\delta_{\textbf{k}\textbf{k}^{\prime}}\,.\end{aligned}\right.$ (S20)
So, from Eq. (S16), we get
$\epsilon_{\gamma\alpha\beta}$ ${\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P\}=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\langle\partial_{k_{x}}u_{n\textbf{k}}|H_{\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ (S21)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ.}}\sum_{m\in{\rm emp.}}\sum_{s\,l}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})E_{m\textbf{k}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle u_{m\textbf{k}}|\partial_{k_{\beta}}u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}.$
Similarly, Eq. (S17) becomes
$\epsilon_{\gamma\alpha\beta}$ ${\rm Tr}\{Qr_{\alpha}PHPr_{\beta}Q\}=-\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}$ (S22)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\sum_{m,m^{\prime}\in{\rm emp.}}\sum_{s\,l}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}.$
@@
$M_{z}$ $=\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ.}}\sum_{m\in{\rm emp.}}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})E_{m\textbf{k}}\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{m\textbf{k}}|u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$ (S23)
$-\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\sum_{m,m^{\prime}\in{\rm emp.}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$
@@
where the layer-resolved orbital magnetization of the $l$-th layer $M_{z}^{(l)}$ is defined by summing all the degrees of freedom except layer $l$. As discussed before, when considering Chern insulators having surface contribution, one needs to replace $H\rightarrow H-\mu$ and $E_{n\textbf{k}}\rightarrow E_{n\textbf{k}}-\mu$ in the previous formula.
@@ I.3 Layer Chern number
$C_{\gamma}(\textbf{r})=\frac{i}{4\pi}\epsilon_{\gamma\alpha\beta}\langle\textbf{r}|Pr_{\alpha}Qr_{\beta}P-Qr_{\alpha}Pr_{\beta}Q|\textbf{r}\rangle.$ (S24)
@@
$C_{z}$ $=-\frac{1}{4\pi}\epsilon_{z\alpha\beta}\,{\rm Im\,Tr}\{Pr_{\alpha}Qr_{\beta}P-Qr_{\alpha}Pr_{\beta}Q\}$ (S25)
@@
II Layer-resolved orbital magnetization under finite vertical electric field
We show in Fig. S1 that $M_{z}^{(l)}$ varies linearly with weak electric field for both types of stacking, independent of number of layers. Here, we calculate $M_{z}^{(l)}$ using the definition Eq. (S23).
Figure S1: Layer-resolved $M_{z}^{(l)}$ of the slab under several $\mathcal{E}_{z}$ for different stacking ways and total layer numbers: (a) AA stacking, 5 layers; (b) AA stacking, 6 layers; (c) AB stacking, 5 layers; (d) AB stacking, 6 layers.
III Perturbation theory for layer-resolved orbital magnetoelectric response
When a vertical electric field is applied to a slab, our perturbed Hamiltonian reads: $H=H_{0}+e\mathcal{E}z$. Suppose the electric field is so weak such that bulk gap is not closed and the potential drop between the top and bottom layer is the smallest energy scale in the system. The idea is to express Eq. (S23) of perturbed Hamiltonian $H$ using the information of $H_{0}$. By singling out the terms proportional to $\mathcal{E}$, the layer-resolved orbital magnetoelectric response $\alpha_{zz}^{(l)}$ is then derived by $\partial M_{z}^{(l)}/\partial\mathcal{E}$. After gathering all the first-order corrections (including energy, wavefunction and chemical potential with particle number fixed), a semi-analytic formula for $\alpha_{zz}^{(l)}$ can be derived. We say “semi-analytic” since the determination of the chemical potential still requires solving numerically the system with open boundary condition, which is unavoidable due to the presence of gapless surface states in 3D Chern insulators. The detailed derivation below can be skipped if the readers are not interested in long but direct algebra. The derived formula is benchmarked to be correct as evidenced by finite-difference method as shown in Table I  in the main text and also Table S1.
Let us first focus on the first-order correction in $\mathcal{E}$ to all terms in Eq. (S23). First, the correction in wavefunction reads
$|u_{n\textbf{k}}\rangle$ $=|u_{n\textbf{k}}^{(0)}\rangle+\sum_{n^{\prime}\neq n}\frac{\frac{1}{N_{\rm cell}}\langle u_{n^{\prime}\textbf{k}}^{(0)}|e\mathcal{E}z|u_{n\textbf{k}}^{(0)}\rangle}{E_{n\textbf{k}}^{(0)}-E_{n^{\prime}\textbf{k}}^{(0)}}|u_{n^{\prime}\textbf{k}}^{(0)}\rangle$ (S26)
@@ III Perturbation theory for layer-resolved orbital magnetoelectric response
$C_{sl,n}(\textbf{k})$ $\equiv\langle sl,\textbf{k}|u_{n\textbf{k}}\rangle$ (S27)
@@
$|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle=-iA_{n,\alpha}(\textbf{k})|u_{n\textbf{k}}\rangle+\sum_{m\neq n}\frac{\frac{1}{N_{\rm cell}}|u_{m\textbf{k}}\rangle\langle u_{m\textbf{k}}|}{E_{n\textbf{k}}-E_{m\textbf{k}}}\,(\partial_{k_{\alpha}}H_{\textbf{k}})|u_{n\textbf{k}}\rangle,$ (S28)
with Berry connection $A_{n,\alpha}(\textbf{k})=i\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle$. Note from Eq. (S23) that we only need to calculate $\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ with $m\neq n$ so we can omit the first term in Eq. (S28). We also note that $\partial_{k_{\alpha}}H_{\textbf{k}}=\partial_{k_{\alpha}}H_{\textbf{k}}^{(0)}$ with $H_{\textbf{k}}^{(0)}=e^{-i\textbf{k}\cdot\textbf{r}}H_{0}\,e^{i\textbf{k}\cdot\textbf{r}}$. Therefore we need to single out the first-order correction to $|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle$ by expanding all other terms. The simplest one would be $E_{n\textbf{k}}$
$E_{n\textbf{k}}$ $=E_{n\textbf{k}}^{(0)}+\langle u_{n\textbf{k}}^{(0)}|e\mathcal{E}z|u_{n\textbf{k}}^{(0)}\rangle$ (S29)
@@
Let’s now expand the second term in Eq. (S28). Since
@@
$\frac{1}{N_{\rm cell}}\sum_{m\neq n}\frac{|u_{m\textbf{k}}\rangle\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}H_{\textbf{k}}|u_{n\textbf{k}}\rangle}{E_{n\textbf{k}}-E_{m\textbf{k}}}$ (S30)
@@
With all the ingredients derived above, we can expand Eq. (S23) to the first-order of $\mathcal{E}$ field. Typically, we need to treat terms like:
@@
((*)( ∗ ) $=C_{sl,n}^{(0)}(\textbf{k})C_{sl,n^{\prime}}^{(0)*}(\textbf{k})E_{m\textbf{k}}^{(0)}\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{m\textbf{k}}^{(0)}|u_{n^{\prime}\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ (S31)
@@
$\langle u_{n\textbf{k}}^{(0)}|u_{m^{\prime}\textbf{k}}^{(1)}\rangle_{\Omega_{0}}$ $=\sum_{s\,l}\sum_{m^{\prime\prime}\neq m^{\prime}}\langle u_{n\textbf{k}}^{(0)}|u_{m^{\prime\prime}\textbf{k}}^{(0)}\rangle_{\Omega_{0}}\frac{C_{sl,m^{\prime\prime}}^{(0)*}(\textbf{k})C_{sl,m^{\prime}}^{(0)}(\textbf{k})}{E_{m^{\prime}\textbf{k}}^{(0)}-E_{m^{\prime\prime}\textbf{k}}^{(0)}}\,e\mathcal{E}ld$ (S35)
@@
$v_{\alpha,nm}(\textbf{k})$ $\equiv\langle u_{n\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}},$ (S36)
@@
$\langle u_{n\textbf{k}}^{(1)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ $=\sum_{n^{\prime}\neq n}\sum_{s\,l}\frac{C_{sl,n^{\prime}}^{(0)}(\textbf{k})C_{sl,n}^{(0)*}(\textbf{k})}{E_{n\textbf{k}}^{(0)}-E_{n^{\prime}\textbf{k}}^{(0)}}\,e\mathcal{E}ld\langle u_{n^{\prime}\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}$ (S37)
@@
$\langle u_{n\textbf{k}}^{(0)}|v_{\alpha}|u_{m\textbf{k}}^{(1)}\rangle_{\Omega_{0}}$ $=\sum_{m^{\prime}\neq m}\sum_{s\,l}\Big{(}\frac{D_{sl,mm^{\prime}}(\textbf{k})}{\Delta_{\textbf{k},mm^{\prime}}}\tilde{\mathcal{E}}\,lv_{\alpha,m^{\prime}n}(\textbf{k})\Big{)}^{*}$ (S38)
@@
So, Eq.(S34) becomes:
$\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(1)}\rangle_{\Omega_{0}}=$ $-\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}^{2}}\sum_{s\,l}\big{(}D_{sl,mm}(\textbf{k})-D_{sl,nn}(\textbf{k})\big{)}\tilde{\mathcal{E}}\,l$ (S39)
@@
Eq.(S33) becomes:
$\langle u_{n\textbf{k}}^{(1)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}=\sum_{n^{\prime}\neq n,m;s\,l}\frac{D_{sl,nn^{\prime}}(\textbf{k})v_{\alpha,n^{\prime}m}(\textbf{k})}{\Delta_{\textbf{k},nn^{\prime}}\Delta_{\textbf{k},mn^{\prime}}}\,\tilde{\mathcal{E}}\,l.$ (S40)
Eq.(S32) becomes:
$\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle_{\Omega_{0}}=\frac{v_{\alpha,nm}(\textbf{k})}{\Delta_{\textbf{k},mn}}.$ (S41)
In Eq. (S27), the first-order term becomes:
$C_{sl,n}^{(1)}(\textbf{k})=\sum_{n^{\prime}\neq n;s^{\prime}l^{\prime}}\frac{D_{s^{\prime}l^{\prime},n^{\prime}n}(\textbf{k})}{\Delta_{\textbf{k},nn^{\prime}}}\,\tilde{\mathcal{E}}\,l^{\prime}C_{sl,n^{\prime}}^{(0)}(\textbf{k}).$ (S42)
@@
So, we decompose Eq. (S31) into four contributions: $(B_{1})$ to $(B_{4})$.
We now explicit the first-order correction in $|u_{n\textbf{k}}\rangle$: Eq.(S39) and Eq.(S40)
$\langle u_{n\textbf{k}}^{(1)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(0)}\rangle+\langle u_{n\textbf{k}}^{(0)}|\partial_{k_{\alpha}}u_{m\textbf{k}}^{(1)}\rangle\equiv-iA_{\alpha,nm}^{(1)}\qquad\qquad\quad({\rm correction\;in\;Berry\;connection})$ (S43)
@@
Now, we are ready for Eq. (S23) of the perturbed system. The first part reads
@@
Let us first look what to change if considering the correction in $\mu$. First, only substitutions $E_{m\textbf{k}}^{(0)}\rightarrow E_{m\textbf{k}}^{(0)}-\mu^{(0)}$ should be made in $(B_{2})$ and $(B_{4})$ because the rest of terms is already of first-order. As for $(B_{3})$, we should take $\mu^{(1)}$ into consideration. Our solution to $\mu^{(1)}$ is to use an auxillary system with OBC, conjugate to the system with PBC. Here, we mean “conjugate” by that both systems have the same level of discretization, in other words the back and forth Fourier transformation is normalized to unity. For example, if we consider a system with PBC of k mesh $100\times 100$ in the Brillouin zone, its conjugate OBC system consists of $100\times 100$ unit cells. For isolated system, a weak external electric field cannot change the particle number of system. So, if the system is half-filled, one determine $\mu^{(1)}$ by
$\mu^{(1)}=\frac{1}{2}(E^{(1)}_{{\rm OBC},v}+E^{(1)}_{{\rm OBC},c})$ (S44)
@@
$\mu^{(1)}=\frac{1}{2}\sum_{\textbf{R},sl}(\tilde{C}_{sl,v}^{(0)\,*}(\textbf{R})\tilde{C}_{sl,v}^{(0)}(\textbf{R})+\tilde{C}_{sl,c}^{(0)\,*}(\textbf{R})\tilde{C}_{sl,c}^{(0)}(\textbf{R}))\,\tilde{\mathcal{E}}l$ (S45)